# Artificial Intelligence NN Neural Network

### v2

@@ 1 Introduction
A novel methodology in this context involves enhancing ML through the use of logical propositions Granmo (2018). This approach enables the representation of a target word by a set of words crucial in shaping its meaning. One key distinction of embeddings generated using logical propositions, as opposed to traditional Deep Learning (DL) methods, is their interpretability. The output, represented by clauses, can be understood and analyzed in line with the original logical framework. In previous work Bhattarai et al. (2024), embeddings were derived for a vector of target words. That method produces a fused output that obscures the individual contributions of each input word, thus forfeiting the interpretability advantage inherent in the original TM algorithm.
An experiment utilizing the One Billion Word dataset Chelba et al. (2013) and the Tsetlin Machine Auto-Encoder (TM-AE) model Bhattarai et al. (2024) revealed that training the model on 100 input target words required approximately 9 hours and 35 minutes. Notably, the embedding generated during each training session is specific to the particular set of input target words and cannot be reused in scenarios involving the addition, removal, or substitution of any element within the input array. Consequently, the time-intensive training process is bound to the initial input string, limiting its reusability in other applications. This scenario highlights the scalability challenges associated with the TM-AE model Bhattarai et al. (2024) when applied to downstream tasks, underscoring the need for further optimization and improvements to enhance its efficiency.
@@
2 Background and Related Work
In the realm of AI applications, NLP has witnessed remarkable advancements over the past decade. Prominent algorithms such as Word2Vec and GloVe have played pivotal roles in this progress Goldberg and Levy (2014); Pennington et al. (2014). Despite the promising results achieved by these embedding algorithms, there remains a pressing need for further improvements. For instance, FastText focuses on subword information, thereby enhancing its capability to handle rare words and misspellings Bojanowski et al. (2017), while ELMo generates contextualized word embeddings by considering the entire sentence Peters et al. (2018).
A novel approach to representing ML through a logical structure is the TM, which emphasizes transparency and the ability to provide justifications for outcomes. TM have seen significant advancements through various contributions. Maheshwari et al. (2023) introduced REDRESS, a method for generating compressed models tailored for edge inference using TM. Sharma et al. (2023) enhanced the robustness and pattern recognition capabilities of TM through the Drop Clause technique. Seraj et al. (2022) explored the application of TM to solve contextual bandit problems. Yadav et al. (2021) demonstrated the human-level interpretability of TM in aspect-based sentiment analysis. Abeyrathna et al. (2021) developed a massively parallel and asynchronous TM architecture for efficient scaling. Additionally, Abeyrathna et al. (2023) proposed a method to build concise logical patterns by constraining the clause size in TM.
In the field of NLP, several models have leveraged TM for various applications. For instance, in the work Yadav et al. (2022), robust and interpretable text classification models were developed by learning logical AND rules with negation. Another study Saha et al. (2023) used TM to discover interpretable rules for tasks such as sentiment analysis, semantic relation categorization, and dialogue-based entity identification. The work Zhang et al. (2023) applied TM in sentiment analysis and spam review detection for Chinese text, aiming to strike a balance between interpretability and accuracy when compared to DL models. Additionally, a TM-AE was designed to identify word embeddings Bhattarai et al. (2024) using the Coalesced Tsetlin Machine (CoTM) structure Glimsdal and Granmo (2021), which incorporates voting on a set of outputs with weighted clauses representing their contributions to the algorithm. That embedding method has demonstrated superior performance compared to DL alternatives.
3 Methodology
@@
3.1 Coalesced Tsetlin Machine
@@
To explain the formulation of the input $X$, let us look at an example. We assume that the vocabulary contains eight words: $word_{1}$, $word_{2}$, …, $word_{8}$. Suppose we have two documents: $document_{1}$, which contains the words [$word_{3},word_{2},word_{4}$] and $document_{2}$, which contains the words [$word_{6},word_{3}$] and we want to form $X$ for the target word $word_{3}$. For the first part of $X$, we select the features corresponding to the words in $document_{1}$ and $document_{2}$ and set their values to $1$. The rest of the features are set to 00. Thus, we have: $X_{original}=[0,1,1,1,0,1,0,0],$ where $X_{original}$ is a binary vector of length 8. For the second part of $X$, we take the negation of $X_{original}$. This means setting the value of each feature to the opposite of its value in $X_{original}$. Thus, we have: $X_{negation}=[1,0,0,0,1,0,1,1].$ Finally, we concatenate $X_{original}$ and $X_{negation}$ to form $X$, as $X=Concat[X_{original},X_{negation}],$ namely, $X=[0,1,1,1,0,1,0,0,1,0,0,0,1,0,1,1].$ This $X$ represents the active literals required to train the TM for the target $word3$, and it has a length of 16 literals. The two phases used in this work are distinguished by the approach employed in forming $X$, to be explained in the next section.
@@
The process of evaluating the clauses begins with the input $X$, where the literals of the input are matched with the propositional logic represented by the conjunctive clauses. These clauses are different forms of description of the target word, composed of literals. The literals memorized in a specific clause are determined during the training process. More specifically, an election process occurs throughout the training period of the model to update the depth of memorizing or forgetting the literals in memory and thus include or exclude them from the clause. Each literal corresponds to a Tsetlin automaton (TA), which decides to memorize or forget the literal based on the reward/penalty dynamics that it receives. In more detail, each TA is assigned to a specific literal in $X$ and evaluates its state across a range of $2N$ levels, where level 1 corresponds to the most extensive forgetting and level $2N$ represents the highest degree of memorization. Fig. 1 illustrates the operation of the TM, highlighting how the TA collaborate to evaluate the input $X$, thereby forming the clause. The input $X$ consists of $V$ features, resulting in a total of $2V$ literals. In the figure, both $x_{1}$ and $\neg x_{V}$ have successfully progressed beyond the $N$ state, thereby contributing to the formation of the clause, represented by the blue area. For more details on the clause formation, see Granmo (2018).
@@
Each clause is updated based on the input $X$ through three types of updates Granmo (2018):
@@
3.2 Tsetlin Machine Autoencoder
In the architecture proposed in Bhattarai et al. (2024), the input $X$ is constructed by randomly shuffling a set of target words. The process for preparing the input $X$ is outlined in Algorithm 2 in the appendix. This algorithm involves combining input documents, where the model receives the documents in their vectorized form, with the vocabulary serving as a reference to track the position of each word within the encoding throughout the training process.
@@ 3.2 Tsetlin Machine Autoencoder
In practical NLP applications, the embedding method in Bhattarai et al. (2024) is not scalable or reusable because the output cannot be effectively leveraged. For example, in LLMs such as GPT-3, one of the initial steps after receiving the input vector is to generate embeddings for each word. These embeddings include context information extracted from documents that the model was trained on, and are stored as dense vector representations in a high-dimensional space. Unlike GPT-3, which retains these embeddings for efficient reuse across tasks, the TM-AE model Bhattarai et al. (2024) generates embeddings specific to the input target words and can not be used for different input combinations. In this work, the aforementioned problems were identified, and a TM-AE was employed in two phases, each with a distinct structure to construct the input $X$ and then trained using the CoTM structure to represent embeddings for a vector of target words. Fig. 3 illustrates the two phases used for this application.
3.3 Phase 1
@@
Documents picked for $\text{Documents picked for }word_{k}=\min(a,|D_{word_{k}}|)$ (1)
@@
Words extracted for $\text{Words extracted for }word_{k}=\sum_{d_{i}\in D_{\text{picked}}}\text{Words}(d_{i})$ (2)
@@ 3.4 Phase 2
Overall, when comparing the computational complexity of the previous Bhattarai et al. (2024) and proposed methods, the following observations can be made:
@@ 4.1 Similarity Evaluation
[…]
Table 1: Comparison of Spearman (S), Kendall (K), and Cosine (C) similarity measures across different datasets and models.
Table 1 shows the evaluation results compared to related algorithms. Despite the potential for scaling and transparency in extracting knowledge for each word in the first phase, our results indicate that the old method is generally more accurate when evaluated using pre-prepared datasets. There are two main reasons for this discrepancy. Firstly, the old method extracts common context information between words by training the TM to fuse and blend all the information in the input samples and find the commonality with respect to the input vector. The training focuses on extracting context information for the input vector as a whole rather than general knowledge for each word in the input vector. In contrast, the first phase of our proposed approach trains the TM in an unsupervised way with the goal of extracting context information for the target word. This may bias the output based on the type of data available in the database and the size of the database. Secondly, the size of the database is crucial. In our proposed approach, the first phase involves training with a specific hyperparameter to extract knowledge that determines finding the embedding for the input vector in the second phase. Even if we were to expand and increase the values of the training for the second phase, it would still be limited by the knowledge extracted from the first phase.
@@
Notably, in terms of Cosine similarity, the two-phase TM-AE achieved 0.91 on the RG65 dataset, surpassing all Word2Vec (0.84), FastText (0.69) and the TM-AE (0.87). This indicates that the two-phase approach effectively captures semantic relationships when similarity is assessed through vector alignment. However, its performance in Spearman and Kendall correlations is comparatively weaker. For instance, in the WS-353 dataset, it scores 0.41 and 0.29, respectively, which is lower than Word2Vec’s scores of 0.58 and 0.41. This suggests that while the two-phase model produces closely aligned embeddings, it may not accurately reflect the rank order of word pairs as judged by humans. Also, when comparing the similarity results between Word2Vec and FastText—which is expected to be an improved version of Word2Vec Bojanowski et al. (2017)—it was observed that FastText produced weaker results in this specific application. This suggests that while the adopted classification is critical for evaluating our model’s performance, it heavily relies on human-created databases. These databases, while generally weighted to measure similarity between words, may not be entirely objective for this application. Furthermore, the two-phase model exhibits variable performance across datasets. For example, it achieves a high Cosine similarity of 0.97 in the MTURK287 dataset but experiences a decline in Spearman correlation to 0.40. Overall, while the two-phase TM-AE shows promise in generating aligned embeddings, further improvements should focus on optimizing the knowledge extraction phase to ensure better alignment with human semantic evaluations across diverse datasets.
@@
Table 2: Comparison of accuracy using different embedding sources and classifier: LR (Logistic Regression), NB (Naive Bayes), RF (Random Forest), SVM (Support Vector Machine), MLP (Multi-layer Perceptron), and TM (Tsetlin Machine Classifier)
4.2 Sentiment Analysis
@@
Table 2 presents our experiments’ accuracy results using the IMDB dataset, which consists of 25K training samples and an additional 25K samples for evaluation. The data augmentation experiments were performed using various embedding models, including GloVe, Word2Vec, FastText, ELMo, BERT, and Two-phase TM-AE. The performance was evaluated using several classifiers: Logistic Regression, Naive Bayes, Random Forest, Support Vector Machine, Multi-Layer Perceptron, and Tsetlin Machine.
@@
The results in Table 2 provide a comprehensive comparison of the performance of various embedding models across different classifiers. The classification was performed with default classifier settings across all experiments and for all embedding sources. For the MLP classifier, the random_state parameter was set to 42. Notably, the two-phase TM-AE model shows competitive accuracy, especially with the TM classifier, achieving an accuracy of 0.78, surpassing GloVe (0.60) and FastText (0.74). However, The two-phase TM-AE falls behind models such as ELMo and BERT. ELMo’s dynamic contextualized embeddings adapt to each word’s context in a sentence, unlike the static embeddings used in the two-phase TM-AE approach. ELMo’s approach consistently selects similar or dissimilar words from a limited range of options within individual documents, as its embeddings are generated per document rather than across the entire dataset. This localized embedding strategy impacts the augmentation of the training set, subsequently influencing the classification results. As for BERT, its superior performance is attributed to its bidirectional nature, which allows it to capture context from both directions in a sentence, leading to richer contextualized embeddings. Additionally, BERT’s transformer-based architecture enables it to model long-range dependencies more effectively, further enhancing classification accuracy.
@@
5 Future Work
In this work, we have presented an optimal approach for initiating the development of a practical embedding that can be leveraged in various applications. The application of sentiment analysis through data augmentation (Section 4.2) represents the first instance in which the embedding method proposed in the first phase was utilized. Looking ahead, our future efforts will focus on enhancing the efficiency of the embedding, improving model performance, and addressing the current limitations associated with the slow implementation.

### v3

@@ I Introduction
A novel methodology in this context involves enhancing Machine Learning (ML) through the use of logical propositions [1]. This approach enables the representation of a target word by a set of words crucial in shaping its meaning. One key distinction of embeddings generated using logical propositions, as opposed to traditional Deep Learning (DL) methods, is their interpretability. The output, represented by clauses, can be understood and analyzed in line with the original logical framework. In previous work [2], embeddings were derived for a vector of target words. That method produces a fused output that obscures the individual contributions of each input word, thus forfeiting the interpretability advantage inherent in the original Tsetlin Machine (TM) algorithm.
An experiment utilizing the One Billion Word dataset [3] and the Tsetlin Machine Auto-Encoder (TM-AE) model [2] revealed that training the model on 100 input target words required approximately 9 hours and 35 minutes. Notably, the embedding generated during each training session is specific to the particular set of input target words and cannot be reused in scenarios involving the addition, removal, or substitution of any element within the input array. Consequently, the time-intensive training process is bound to the initial input string, limiting its reusability in other applications. This scenario highlights the scalability challenges associated with the TM-AE model [2] when applied to downstream tasks, underscoring the need for further optimization and improvements to enhance its efficiency.
@@
II Background and Related Work
In the realm of AI applications, Natural Language Processing (NLP) has witnessed remarkable advancements over the past decade. Prominent algorithms such as Word2Vec and GloVe have played pivotal roles in this progress [4, 5]. Despite the promising results achieved by these embedding algorithms, there remains a pressing need for further improvements. For instance, FastText focuses on subword information, thereby enhancing its capability to handle rare words and misspellings [6], while ELMo generates contextualized word embeddings by considering the entire sentence [7].
A novel approach to representing ML through a logical structure is the TM, which emphasizes transparency and the ability to provide justifications for outcomes. TM have seen significant advancements through various contributions. [8] introduced REDRESS, a method for generating compressed models tailored for edge inference using TM. [9] enhanced the robustness and pattern recognition capabilities of TM through the Drop Clause technique. [10] explored the application of TM to solve contextual bandit problems. [11] demonstrated the human-level interpretability of TM in aspect-based sentiment analysis. [12] developed a massively parallel and asynchronous TM architecture for efficient scaling. Additionally, [13] proposed a method to build concise logical patterns by constraining the clause size in TM.
In the field of NLP, several models have leveraged TM for various applications. For instance, in the work [14], robust and interpretable text classification models were developed by learning logical AND rules with negation. Another study [15] used TM to discover interpretable rules for tasks such as sentiment analysis, semantic relation categorization, and dialogue-based entity identification. The work [16] applied TM in sentiment analysis and spam review detection for Chinese text, aiming to strike a balance between interpretability and accuracy when compared to DL models. Additionally, a TM-AE was designed to identify word embeddings [2] using the Coalesced Tsetlin Machine (CoTM) structure [17], which incorporates voting on a set of outputs with weighted clauses representing their contributions to the algorithm. That embedding method has demonstrated superior performance compared to DL alternatives.
III Methodology
@@
III-A Coalesced Tsetlin Machine
@@
To explain the formulation of the input $X$, let us look at an example. We assume that the vocabulary contains eight words: $word_{1}$, $word_{2}$, …, $word_{8}$. Suppose we have two documents: $document_{1}$, which contains the words $word_{3},word_{2},word_{4}$] and $document_{2}$, which contains the words $word_{6},word_{3}$] and we want to form $X$ for the target word $word_{3}$. For the first part of $X$, we select the features corresponding to the words in $document_{1}$ and $document_{2}$ and set their values to $1$. The rest of the features are set to 0. Thus, we have: $X_{original}=[0,1,1,1,0,1,0,0],$ where $X_{original}$ is a binary vector of length 8. For the second part of $X$, we take the negation of $X_{original}$. This means setting the value of each feature to the opposite of its value in $X_{original}$. Thus, we have: $X_{negation}=[1,0,0,0,1,0,1,1].$ Finally, we concatenate $X_{original}$ and $X_{negation}$ to form $X$, as $X=Concat[X_{original},X_{negation}],$ namely, $X=[0,1,1,1,0,1,0,0,1,0,0,0,1,0,1,1].$ This $X$ represents the active literals required to train the TM for the target $word3$, and it has a length of 16 literals. The two phases used in this work are distinguished by the approach employed in forming $X$, to be explained in the next section.
@@
The process of evaluating the clauses begins with the input $X$, where the literals of the input are matched with the propositional logic represented by the conjunctive clauses. These clauses are different forms of description of the target word, composed of literals. The literals memorized in a specific clause are determined during the training process. More specifically, an election process occurs throughout the training period of the model to update the depth of memorizing or forgetting the literals in memory and thus include or exclude them from the clause. Each literal corresponds to a Tsetlin automaton (TA), which decides to memorize or forget the literal based on the reward/penalty dynamics that it receives. In more detail, each TA is assigned to a specific literal in $X$ and evaluates its state across a range of $2N$ levels, where level 1 corresponds to the most extensive forgetting and level $2N$ represents the highest degree of memorization. Fig. 1 illustrates the operation of the TM, highlighting how the TA collaborate to evaluate the input $X$, thereby forming the clause. The input $X$ consists of $V$ features, resulting in a total of $2V$ literals. In the figure, both $x_{1}$ and $\neg x_{V}$ have successfully progressed beyond the $N$ state, thereby contributing to the formation of the clause, represented by the blue area. For more details on the clause formation, see [1].
@@
Each clause is updated based on the input $X$ through three types of updates [1]:
@@
III-B Tsetlin Machine Autoencoder
In the architecture proposed in [2], the input $X$ is constructed by randomly shuffling a set of target words. The process for preparing the input $X$ is outlined in Algorithm 2 in the appendix. This algorithm involves combining input documents, where the model receives the documents in their vectorized form, with the vocabulary serving as a reference to track the position of each word within the encoding throughout the training process.
@@ III-B Tsetlin Machine Autoencoder
In practical NLP applications, the embedding method in [2] is not scalable or reusable because the output cannot be effectively leveraged. For example, in LLMs such as GPT-3, one of the initial steps after receiving the input vector is to generate embeddings for each word. These embeddings include context information extracted from documents that the model was trained on, and are stored as dense vector representations in a high-dimensional space. Unlike GPT-3, which retains these embeddings for efficient reuse across tasks, the TM-AE model [2] generates embeddings specific to the input target words and can not be used for different input combinations. In this work, the aforementioned problems were identified, and a TM-AE was employed in two phases, each with a distinct structure to construct the input $X$ and then trained using the CoTM structure to represent embeddings for a vector of target words. Fig. 3 illustrates the two phases used for this application.
III-C Phase 1
@@
$\text{Documents picked for }word_{k}=\min(a,|D_{word_{k}}|)$ (1)
@@
$\text{Words extracted for }word_{k}=\sum_{d_{i}\in D_{\text{picked}}}\text{Words}(d_{i})$ (2)
@@ III-D Phase 2
Overall, when comparing the computational complexity of the previous [2] and proposed methods, the following observations can be made:
@@ IV-A Similarity Evaluation
Table I: Comparison of Spearman (S), Kendall (K), and Cosine (C) similarity measures across different datasets and models.
@@
Avg. 0.52 0.36 0.87 0.47 0.32 0.69 0.50 0.34 0.87 0.41 0.30 0.93
Table I shows the evaluation results compared to related algorithms. Despite the potential for scaling and transparency in extracting knowledge for each word in the first phase, our results indicate that the old method is generally more accurate when evaluated using pre-prepared datasets. There are two main reasons for this discrepancy. Firstly, the old method extracts common context information between words by training the TM to fuse and blend all the information in the input samples and find the commonality with respect to the input vector. The training focuses on extracting context information for the input vector as a whole rather than general knowledge for each word in the input vector. In contrast, the first phase of our proposed approach trains the TM in an unsupervised way with the goal of extracting context information for the target word. This may bias the output based on the type of data available in the database and the size of the database. Secondly, the size of the database is crucial. In our proposed approach, the first phase involves training with a specific hyperparameter to extract knowledge that determines finding the embedding for the input vector in the second phase. Even if we were to expand and increase the values of the training for the second phase, it would still be limited by the knowledge extracted from the first phase.
@@
Notably, in terms of Cosine similarity, the two-phase TM-AE achieved 0.91 on the RG65 dataset, surpassing all Word2Vec (0.84), FastText (0.69) and the TM-AE (0.87). This indicates that the two-phase approach effectively captures semantic relationships when similarity is assessed through vector alignment. However, its performance in Spearman and Kendall correlations is comparatively weaker. For instance, in the WS-353 dataset, it scores 0.41 and 0.29, respectively, which is lower than Word2Vec’s scores of 0.58 and 0.41. This suggests that while the two-phase model produces closely aligned embeddings, it may not accurately reflect the rank order of word pairs as judged by humans. Also, when comparing the similarity results between Word2Vec and FastText—which is expected to be an improved version of Word2Vec [6]—it was observed that FastText produced weaker results in this specific application. This suggests that while the adopted classification is critical for evaluating our model’s performance, it heavily relies on human-created databases. These databases, while generally weighted to measure similarity between words, may not be entirely objective for this application. Furthermore, the two-phase model exhibits variable performance across datasets. For example, it achieves a high Cosine similarity of 0.97 in the MTURK287 dataset but experiences a decline in Spearman correlation to 0.40. Overall, while the two-phase TM-AE shows promise in generating aligned embeddings, further improvements should focus on optimizing the knowledge extraction phase to ensure better alignment with human semantic evaluations across diverse datasets.
Table II: Comparison of accuracy using different embedding sources and classifier: LR (Logistic Regression), NB (Naive Bayes), RF (Random Forest), SVM (Support Vector Machine), MLP (Multi-layer Perceptron), and TM (Tsetlin Machine Classifier)
@@
IV-B Sentiment Analysis
@@
Table II presents our experiments’ accuracy results using the IMDB dataset, which consists of 25K training samples and an additional 25K samples for evaluation. The data augmentation experiments were performed using various embedding models, including GloVe, Word2Vec, FastText, ELMo, BERT, and Two-phase TM-AE. The performance was evaluated using several classifiers: Logistic Regression, Naive Bayes, Random Forest, Support Vector Machine, Multi-Layer Perceptron, and Tsetlin Machine.
@@
The results in Table II provide a comprehensive comparison of the performance of various embedding models across different classifiers. The classification was performed with default classifier settings across all experiments and for all embedding sources. For the MLP classifier, the random_state parameter was set to 42. Notably, the two-phase TM-AE model shows competitive accuracy, especially with the TM classifier, achieving an accuracy of 0.78, surpassing GloVe (0.60) and FastText (0.74). However, The two-phase TM-AE falls behind models such as ELMo and BERT. ELMo’s dynamic contextualized embeddings adapt to each word’s context in a sentence, unlike the static embeddings used in the two-phase TM-AE approach. ELMo’s approach consistently selects similar or dissimilar words from a limited range of options within individual documents, as its embeddings are generated per document rather than across the entire dataset. This localized embedding strategy impacts the augmentation of the training set, subsequently influencing the classification results. As for BERT, its superior performance is attributed to its bidirectional nature, which allows it to capture context from both directions in a sentence, leading to richer contextualized embeddings. Additionally, BERT’s transformer-based architecture enables it to model long-range dependencies more effectively, further enhancing classification accuracy.
@@
V Future Work
In this work, we have presented an optimal approach for initiating the development of a practical embedding that can be leveraged in various applications. The application of sentiment analysis through data augmentation (Section IV-B) represents the first instance in which the embedding method proposed in the first phase was utilized. Looking ahead, our future efforts will focus on enhancing the efficiency of the embedding, improving model performance, and addressing the current limitations associated with the slow implementation.